The most recent financial statements for Moose Tours, Inc., appear below. Sales for 2016 are projected to grow by 25 percent. Interest expense will remain constant, the tax rate and the dividend payout rate will also remain constant. Costs, other expenses, current assets, fixed assets, and accounts payable increase spontaneously with sales. MOOSE TOURS, INC 2015 Income Statement Sales Costs Other expenses $744,000 579,000 15,000 Earnings before interest and taxes Interest expense $150,000 11,000 Taxable income Taxes (20%) $139,000 27,800 Net income $111,200 Dividends Addition to retained earnings $22,240 88,960 MOOSE TOURS, INC Balance Sheet as of December 31, 2015 Assets Liabilities and Owners Equity Current assets Current liabilities $54,500 Cash Accounts receivable Inventory $ 20,340 Acounts payable 32,660 Notes paya 69,620 13,700 $68,200 $127,000 yable Total Total $122,620 Long-term debt Fixed assets Owners equity Net plant and equipment $420,000 Cmmon stock and paid-in surplus $113,000 Retained earnings 234,420 $347,420 $542,620 Total Total assets $542,620 Total liabilities and owners equity If the firm is operating at full capacity and no new debt or equity is issued, what external financing is needed to support the 25 percent growth rate in sales? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) External financing needed 24345
Answer #1

Calculation of additional financing required:Income Statement $744,000 $579,000 $15,000 150,000 $11,000 $139,000 $27,800 $111,200 $22,240 $88,960 Sales Balance sheet Assets Cash Account Receivables$32,660 Inventories Total current assets $122,620 Fixed assets Total Assets Liabilities and Equit Other Expenses EBIT Interest Taxable Income Tax (20%) Net Income Cash Dividends Added retained earnin $54,500 $13,700 $127,000 $195,200 $113,000 $234,420 Total Liabilities and Equity $542,620 20,340 Accounts Payable Notes Payable 69,620Lonterm Debt Total Liabilities 420,000 Comon Stock 542,620 Retained Earnin 10 15 25% Sales Growth Dividend Payout 16 20%)-D12/D11 Additional fund needed (AFN) equation is as follows 21 23 Where, A-current level of assets AS/s-percentage increase in sales i.e. change in sales divided by current sales Lo current level of Spontaneous liabilities (account payables, accrued liaibilities etc.) S1 new level of sales PM profit margin b retention rate 1-payout rate 25 29 744,000 $542,620 54,500 $0 $111,200 31 Current Net Sales, NSO Total Assets, TAO Account Payable, APO Accrrued Liabilities, ALO Net Income, NiO Retention ratio, RRO 32 35 Next Period Sales Growth, g Next year net sales, NS1 Change in net sales, ANS NS1-NSO 25% 930,000 D31 (1+D38) $186,000D39-D31 41 42 43 Change in Total Assets, ΔΤΑ : (TAO / NSO) *SNS Change in Spontaneous Liability, A(AP AL ((APO+ALO)/NS0)ANS Addition to retained earnings (NS1 (NIO/NSO) *RRO $135,655(D32/D31)*D40 13,625 ((D33+D34)/D31) D40 111,200 (D39 (D35/D31) D36) Additional Fund Needed ΔΙΑ)-Δ (AP+AL)-(addition to Retained earnings) 10,830D42-D43-D44 Hence additional fund needed is $10,830

Formula sheet

A B C D E F G H K
2
3 Income Statement
4 Sales 744000 Balance sheet
5 Costs 579000 Assets Liabilities and Equity
6 Other Expenses 15000 Cash 20340 Accounts Payable 54500
7 EBIT =D4-D5-D6 Account Receivables 32660 Notes Payable 13700
8 Interest 11000 Inventories 69620 Long term Debt 127000
9 Taxable Income =D7-D8 Total current assets =SUM(G6:G8) Total Liabilities =SUM(J6:J8)
10 Tax (20%) =D9*20% Fixed assets 420000 Common Stock 113000
11 Net Income =D9-D10 Total Assets =G9+G10 Retained Earnings 234420
12 Cash Dividends 22240 Total Liabilities and Equity =J9+J10+J11
13 Added retained earnings =D11-D12
14
15 Sales Growth 0.25
16 Dividend Payout =D12/D11 =D12/D11
17
18 Additional fund needed (AFN) equation is as follows:
19 Pict Pict Pict
20
21
22
23 Where,
24 A0 = current level of assets
25 ΔS/S0 = percentage increase in sales i.e. change in sales divided by current sales
26 Lo = current level of Spontaneous liabilities (account payables, accrued liaibilities etc.)
27 S1 = new level of sales
28 PM = profit margin
29 b = retention rate = 1 – payout rate
30
31 Current Net Sales, NS0 =D4
32 Total Assets, TA0 =G11
33 Account Payable, AP0 =J6
34 Accrrued Liabilities, AL0 0
35 Net Income, NI0 =D11
36 Retention ratio, RR0 =(1-D16)
37
38 Next Period Sales Growth, g =D15
39 Next year net sales, NS1 =D31*(1+D38) =D31*(1+D38)
40 Change in net sales, ∆NS = NS1-NS0 =D39-D31 =D39-D31
41
42 Change in Total Assets, ΔTA = (TA0 / NS0) * ∆NS =(D32/D31)*D40 =(D32/D31)*D40
43 Change in Spontaneous Liability, Δ(AP + AL)= ((AP0 + AL0)/ NS0) * ∆NS   =((D33+D34)/D31)*D40 =((D33+D34)/D31)*D40
44 Addition to retained earnings = (NS1 * (NI0 / NS0) * RR0) =(D39*(D35/D31)*D36) =(D39*(D35/D31)*D36)
45
46 Additional Fund Needed =(ΔTA) - Δ (AP+AL) - (addition to Retained earnings)
47 =D42-D43-D44 =D42-D43-D44
48
49 Hence additional fund needed is =D47
50